# Exploring Multiple Strategies to Improve Multilingual Coreference Resolution in CorefUD

### v1

@@ 4.6 Overlapping Segments
We propose using segment overlapping with a cluster merging algorithm to address this issue. The cluster merging algorithm is straightforward: We iterate over all mentions in the new part and if it is present in a cluster in the previous part we simply make a union between its clusters from both parts. We recommend using maximal segment overlap. For instance, if a document has 6 segments and the maximum number of segments for a single example is 4, then we split it into 3 examples overlapping with 3 segments. The example is shown in Figure 5. We further extend this method by only using clusters in the continuing examples where at least one mention is present in new segments because if there is no such mention then all the mentions were seen already in the previous part with a longer left context.
@@ 5.2 Baseline and Extensions
We began by training a monolingual model for each language. For these experiments, we use XLM-Roberta-large and specific monolingual models for each language. Monolingual models for individual languages are listed in Table 5. Then we trained a joined cross-lingual model for all the datasets which we fine-tune for each dataset afterwards.
To evaluate the effect of each extension, we first run all its variants on a basic XLM-R-large model and measure the performance for each language (Sections 6.1.1 and 6.1.2). Then we compare each extension in its best configuration to the baseline joined model in Section 6.1.3 (since joined models achieve the best results from our three baseline models). In the last step, we evaluate the effects of Overlapping segments only on the best model configuration for each language (because it is orthogonal to all other extensions).
@@ 5.2.1 Monolingual Models
We try a specific monolingual model for each language. We use a large variant where it is available. For Czech we use Czert [26], for English Roberta-large [27], gbert-large [28]. Other languages these models [29, 30, 31, 32, 33, 34, 35, 36]. The complete list can be found in Table 5.
@@ 5.3 Zero-shot Cross-lingual Experiments
We also conducted zero-shot cross-lingual experiments in two variants; dataset zero shot, and language zero shot. In dataset zero shot we train the model on all the datasets except one and then we evaluate the model on the excluded dataset. However, this approach is not a true cross-lingual zero-shot test, as multiple datasets exist for some languages, meaning the model was still exposed to the language of the evaluation data during training.
@@ 6 Results & Discussion
Table 5 presents the results of baseline models on development data parts. The Monoling column shows the result of the monolingual model specific for each language. The XLM-R column presents results of XLM Roberta large trained separately for each dataset. The Joined column corresponds to the joined model described in Section 4.1. Several key observations can be drawn from these results. Most notably, monolingual models tend to outperform the cross-lingual XLM-Roberta model, especially for the Lithuanian and Czech PCEDT datasets, where the monolingual models—even though they are smaller—exhibit significantly better performance. The reason is probably in the difference of these two datasets from the other ones. We can see that for these two datasets, joined training does not help as much as for other datasets. For the rest of the datasets, joined training surpasses both models (monolingual and multilingual) trained from scratch.
@@ 6.1.1 Span2Head
Table 6 shows the performance of different variants of span2head model. None column shows the results of the baseline XLM-R-large model without any extensions. The Multi column reflects the results of the multi-class relative position classification (the first method described in Section 4.3). The binary column represents the second method (binary classification of span-headword candidate pair). From the results, we can make several conclusions:
@@
    Span2head model slightly helps
@@
    Binary model achieves better results, than the multi-class model
@@ 6.1.2 Singletons
Table 7 shows the results of all proposed variants of the singletons extension. Again, None column represents the baseline xlm-r-large trained from scratch for each dataset. Dummy column shows the results of plain dummy antecedent for singletons. mask extends the previous method binary score masking for singletons. This means that the method does not use the binary score for similarities to singleton dummy antecedent. separate uses separate FFNN for singletons score prediction (different from the one used for standard mentions). the last method mentions represents the method described in section 4.5.2. Additionally, the Table shows the number of parameters added by each variant of the model. The last column shows the percentage of singleton entities in each dataset (datasets with few singletons do not need any singleton model).
@@
From the table, we can see that on average the best results were achieved with the mention model. mask model has very poor results, but it makes sense because the number of specific parameters for singletons modeling is very low (3 092 params is used as embedding for dummy antecedent) and if we discard the binary score, we reduce the learning power of the model even more. Dummy model does not help for any dataset statistically significantly. Separate and Mentions models are significantly better for Spanish (Mentions even for Catalan) which has almost no singletons. One reasonable explanation is that the baseline model is under-parametrized for this dataset. We evaluated this hypothesis by increasing the model size, but the improvement was not statistically significant. In the following chapter, we evaluate singletons on the model with joined pretraining, we compute confidence intervals and in Table 13 we can see that the improvemnt is below statistical significance (on $95\%$ significance level). Another possibility is that the model learns to classify properly most of the 3.3% singletons and this leads to a slight improvement of the overall score.
@@ 6.1.3 Best Combinations
Table 8 presents the performance of all the extensions proposed. The Table also shows the 95% confidence intervals. Span2head and Only head extend Trees (both also uses syntax tree representations). We can see, that tree representations alone improve the results only for ParCor corpora. Span2Head model helps significantly only for English-GUM and Hungarian-korkor corpora but it is not better than only heads model. Singletons modeling helps for the most of the datasets which have singletons annotated. The exceptions are in Norwegian Bokmaal and German Potsdam dataset. Another interesting point is that for German ParCor dataset singletons modeling helps significantly even though de-parcor itself does not contains singletons. One possible explanation is in the joined model. Adding the singletons from other languages to the training data can help to predict the mentions in German ParCor dataset. Results in Table 7 supports this theory, because we can see that when we train the model from scratch, singletons modeling does not help. Only heads model helps for almost all the dataset since it reduces the mention space and improves precision significantly. The extensions selected for the best model are highlighted in green color and the last column shows the results of the best combination. The experiments were performed again so even for the datasets where a single extension was selected the results are slightly different but the confidence intervals match.
Table 8: Performance gains over best base model with proposed extensions together with 95% confidence intervals. The models which surpass the baseline significantly are bold. The results which are better only on lower confidence level are underlined. Variants selected for the best models are green.
@@ 6.1.4 Overlapping Segments
Table 9 shows the evaluation of longer context implementation through overlapping segments. We use a maximum length of 8 segments and we evaluate minimum and maximum overlap (1, and 7 respectively), We also try variants with filtering already-seen mentions as described in section 4.6. For most of the datasets the results do not change at all so additionally, we run the same experiment with just 4 segments (Table 10) and we count basic statistics of long coreference relations (Table 11). Table 11 presents three types of information:
@@
    column nearest coref cross N is the percentage of mentions for which the nearest antecedent is already in a different segments block so all the antecedents are in the different block and the antecedent cannot be predicted with the standard model.
@@
    columnsements over N shows the percentage of segments, which are after the first N-segments block. Only those segments are affected by prediction with segment overlapping.
We can see that a significant number of log-distant coreferences is in French Turkish and Russian datasets. Some of them are also in both Czech datasets, Catalan and Polish. By far the most long coreferences are in the French dataset by from Table 9 we can see that segments overlapping decreases the performance. We investigated the dataset closely and found out that in this dataset multiple documents are merged into a single document quite often. When this happens, cross coreferences between those concatenated documents are not annotated so the model which can predict long-distant coreferences is penalized. To evaluate the models correctly we add the variant where we split the documents in French dev data manually (row fr_split). After splitting results with segments overlapping improves significantly but they are still not as high as we expected in comparison with Russian where the overlap helps significantly. To evaluate this a bit further we computed also the recall for cross-segment coreference links. The recall is 72% for Russian and only 44% for French. By comparing these datasets manually we conclude that the coreference chains are very long in the French dataset and it is probably too hard to predict them correctly from the antecedents (one mistake in the chain might have a huge effect on the overall performance). For the Turkish dataset, it is hard to make any significant improvement due to the low quality of the dataset.
Table 9: Results of the best model configurations with different approaches to merge segments during prediction step. Maximum 8 segments
@@
Table 10: Results of the best model configurations with different approaches to merge segments during prediction step. Maximum 4 segments
@@
Table 11: Statistics of long documents and distant coreference relations accross the datasets.
@@ 6.2 Cross-lingual Transfer
Table 12: Zero-shot cross-lingual transfer results
@@
Table 12 shows the results of cross-lingual transfer evaluation. We can see that for most of the datasets fine-tuning the model on the specific data helps but the difference is not so significant. For several datasets, the results are even comparable. For small datasets, the fine-tuned models achieve much better results than the joined model. Note that we do not use any weighing of particular datasets in the joined collection so small datasets have a marginal effect on the training. The results of dataset zero-shot and language zero-shot are mostly very predictable. For the languages that have more datasets in CorefUD collection, the drop of zero-shot compared to the joined model was smaller than for the rest of the datasets. When we discard all the datasets for each language, the results are very similar for most of the datasets. There are several exceptions that should be analyzed more. The strangest results are on the Turkish dataset. Here the zero-shot model achieves the best performance. Together with the fact that the numbers are very low for this dataset, it suggests that there is some noise in the dataset annotation. If we train the model on the training part of the Turkish dataset the noise is present on two sides (training and testing). If we train a zero-shot model we eliminate one part.
@@ 6.3 Final Results
Table 13 presents the final results of the best-performing models on a test set. For a comparison, we use the best model of CRAC 2023 Multilingual coreference resolution shared task from [37] but not the best submission. The best model is much larger than ours and it uses ensembling so as the main model for comparison we use their RemBERT version without ensembling which has comparable size to our model.
From the Table, we can see that our model outperformed CorPipe on most of the datasets. For some datasets (mostly the ones without singletons annotated) it outperformed even the large CorPipe model which has approximately 3 times more trainable parameters. The only dataset where CorPipe outperformed our model by a large margin is the Polish one which we consider an anomaly because for this dataset, CorPipe results on test set are much better than for the dev set. On dev set we achieve similar results. The same happened in the opposite way for Russian dataset. Surprisingly, we even outperformed the large CorPipe model in average score, but this is caused mainly by by a large margin on Turkish and German-parcor
@@ 7 Conclusion
In this paper, we explored and evaluated various approaches to multilingual coreference resolution using the CorefUD 1.1 dataset. Our experiments revealed that monolingual models typically outperform cross-lingual models, especially for languages with datasets that are distinct in their characteristics. However, joint training across languages can still provide benefits for the most of the datasets where sufficient cross-linguistic similarities exist.
We proposed several extensions to enhance the baseline models, including cross-lingual training, Span2Head modeling, syntactic information integration, headword mention representation, and long-context prediction. Among these, the heads-only model and singleton modeling showed the most consistent improvements across different datasets, demonstrating the importance of targeted adaptations for coreference resolution tasks. For several datasets, long-context prediction bring also significant improvement. Additionally, our zero-shot cross-lingual experiments provided insights into the challenges and opportunities of cross-lingual transfer, with the Turkish dataset results highlighting potential issues related to noise in the training data.
@@ Acknowledgement
Computational resources were supplied by the project “e-Infrastruktura CZ” (e-INFRA LM2018140) provided within the program Projects of Large Research, Development and Innovations Infrastructures. This work has been supported by Grant No. SGS-2022-016 Ad- vanced methods of data processing and analysis.

### v2

@@ 4.6 Overlapping Segments
We propose using segment overlapping with a cluster merging algorithm to address this issue. The cluster merging algorithm is straightforward: We iterate over all mentions in the new part and if it is present in a cluster in the previous part we simply make a union between its clusters from both parts. We experiment with minimal (just one overlapping segment) and maximal (each continuing example has just one new segment the rest is overlapping) segment overlap. For instance, if a document has 6 segments and the maximum number of segments for a single example is 4, then we split the document into 3 examples overlapping with 3 segments. The example is shown in Figure 5. We further extend this method by only using clusters in the continuing examples where at least one mention is present in new segments because if there is no such mention then all the mentions were seen already in the previous part with a longer left context.
@@ 5.2 Baseline and Extensions
We begin by training a monolingual model for each language. For these experiments, we use XLM-Roberta-large and specific monolingual models for each language. Monolingual models for individual languages are listed in Table 5. Then we train a joined cross-lingual model for all the datasets which we fine-tune for each dataset afterwards.
To evaluate the effect of each extension, we first run all its variants described in 4 on the basic XLM-R-large model and measure the performance for each language (Sections 6.1.1 and 6.1.2). Then we compare each extension in its best configuration to the baseline joined model in Section 6.1.3 (since joined models achieve the best results from our three baseline models). In the last step, we evaluate the effects of Overlapping segments only on the best model configuration for each language (because it is orthogonal to all other extensions).
@@ 5.2.1 Monolingual Models
We evaluate a specific monolingual model for each language. We use a large variant where it is available. For Czech we use Czert [26], for English Roberta-large [27], gbert-large [28]. Other languages these models [29, 30, 31, 32, 33, 34, 35, 36]. The complete list can be found in Table 5.
@@ 5.3 Zero-shot Cross-lingual Experiments
We also conduct zero-shot cross-lingual experiments in two variants; dataset zero shot, and language zero shot. In dataset zero shot we train the model on all the datasets except one and then we evaluate the model on the excluded dataset. However, this approach is not a true cross-lingual zero-shot test, as multiple datasets exist for some languages, meaning the model was still exposed to the language of the evaluation data during training.
@@ 6 Results & Discussion
Table 5 presents the results of baseline models on development data parts. The Monoling column shows the result of the monolingual model specific to each language. The XLM-R column presents results of XLM Roberta large trained separately for each dataset. The Joined column corresponds to the joined model described in Section 4.1. Several key observations can be drawn from these results. Most notably, monolingual models tend to outperform the cross-lingual XLM-Roberta model, especially for the Lithuanian and Czech PCEDT datasets, where the monolingual models—even though they are smaller—exhibit significantly better performance. The reason is probably in the difference of these two datasets from the other ones. We can see that for these two datasets, joined training does not help as much as for other datasets. For the rest of the datasets, joined training surpasses both other models (monolingual and multilingual trained from scratch).
@@ 6.1.1 Span2Head
Table 6 shows the performance of different variants of span2head model. None column shows the results of the baseline XLM-R-large model without any extensions. The Multi column reflects the results of the multi-class relative position classification (the first method described in Section 4.3). The binary column represents the second method (binary classification of span-headword candidate pair). We added Heads column with heads-only mention representation to show if it is better to use the span2head model or heads only model (since they are mutually exclusive). From the results, we can make several conclusions:
@@
    Span2head model slightly helps.
@@
    Binary model achieves better results, than the multi-class model.
@@ 6.1.2 Singletons
Table 7 shows the results of all proposed variants of the singletons extension. Again, None column represents the baseline xlm-r-large trained from scratch for each dataset. Dummy column shows the results of plain dummy antecedent for singletons. mask extends the previous method binary score masking for singletons. This means that the method does not use the binary score for similarities to the singleton dummy antecedent. separate uses separate FFNN for singletons score prediction (different from the one used for standard mentions). The last method mentions represents the method described in section 4.5.2. Additionally, the Table shows the number of parameters added by each variant of the model. The last column shows the percentage of singleton entities in each dataset (datasets with few singletons do not need any singleton model).
@@
From the table, we can see that on average the best results were achieved with the mention model. Mask model has very poor results, but it makes sense because the number of specific parameters for singletons modeling is very low (3 092 params is used as embedding for dummy antecedent) and if we discard the binary score, we reduce the learning power of the model even more. Dummy model does not help for any dataset statistically significantly. Separate and Mentions models are significantly better for Spanish (Mentions even for Catalan) which has almost no singletons. One reasonable explanation is that the baseline model is under-parametrized for this dataset. We evaluated this hypothesis by increasing the model size, but the improvement was not statistically significant. In the following chapter, we evaluate singletons on the model with joined pretraining, we compute confidence intervals, and in Table 13 we can see that the improvement is below statistical significance (on $95\%$ significance level). So we believe that these strange results in Table 7 are caused by random chance.
@@ 6.1.3 Best Combinations
Table 8 presents the performance of all the extensions proposed. The table also shows the 95% confidence intervals. Span2head and Only head extend Trees (both also uses syntax tree representations). We can see, that tree representations alone improve the results only for ParCor corpora. Span2Head model helps significantly only for English-GUM and Hungarian-korkor corpora but it is not better than only heads model. Singletons modeling helps for most of the datasets that have singletons annotated. The exceptions are in Norwegian Bokmaal and German Potsdam datasets. Another interesting point is that for German ParCor dataset singletons modeling helps significantly even though de-parcor itself does not contain singletons. One possible explanation is in the joined model. Adding the singletons from other languages to the training data can help to predict the mentions in German ParCor dataset. Results in Table 7 support this theory because we can see that when we train the model from scratch, singletons modeling does not help. Only heads model helps for almost all the datasets since it reduces the mention space and improves precision significantly. The extensions selected for the best model are highlighted in green color and the last column shows the results of the best combination. Please note that the results may show slight inconsistencies, even for datasets where a single extension was selected (e.g., ’Only heads’ and ’Best Comb.’ for Hungarian). This is because the experiments were rerun, averaging results across different runs with the same configuration. However, these differences are within the confidence intervals, ensuring the results remain reliable.
Table 8: Performance gains over the best base model with proposed extensions together with 95% confidence intervals. The models that surpass the baseline significantly are bold. The results which are better only on lower confidence level are underlined. Variants selected for the best models are green.
@@ 6.1.4 Overlapping Segments
Table 9 evaluates longer context implementation through overlapping segments. We use a maximum length of 8 segments666since we use this maximum length in most of our other experiments, and we evaluate minimum and maximum overlap (1 and 7, respectively). We also try variants with filtering already-seen mentions as described in section 4.6 (columns min_filter and max_filter). For most of the datasets, the results do not change at all, so additionally, we run the same experiment with just 4 segments (Table 10)777which is the minimum number of segments we used in our models, and we count basic statistics of long coreference relations (Table 11). Table 11 presents three types of information:
@@
    column nearest coref cross N is the percentage of mentions for which the nearest antecedent is already in a different segments block, so all the antecedents are in the different block and the antecedent cannot be predicted with the standard model.
@@
    column segments over N shows the percentage of segments that are after the first N-segments block. Only those segments beyond the first N-segment block are affected by predictions involving segment overlapping, as the remaining segments fit within the model’s capacity and are processed in a single example.
We can see that a significant number of log-distant coreferences are in French, Turkish, and Russian datasets. Some of them are also in both Czech datasets, Catalan and Polish. By far the most long coreferences are in the French dataset. From Table 9, we can see that segments overlapping decreases the performance. We investigated the dataset closely and found out that in this dataset multiple documents are merged into a single document quite often. When this happens, cross coreferences between those concatenated documents are not annotated so the model which can predict long-distant coreferences is penalized. To evaluate the models correctly, we add the variant where we split the documents in French dev data manually (row fr_split). After splitting, results with segments overlapping improve significantly, but they are still not as high as we expected in comparison with Russian, where the overlap helps significantly. To evaluate this a bit further, we also computed the recall for cross-segment coreference links. The recall is 72% for Russian and only 44% for French. By comparing these datasets manually, we conclude that the coreference chains are very long in the French dataset, and it is probably too hard to predict them correctly from the antecedents (one mistake in the chain might have a huge effect on the overall performance). For the Turkish dataset, it is hard to make any significant improvement due to its low quality.
Table 9: Results of the best model configurations with different approaches to merge segments during the prediction step. Maximum 8 segments
@@
avg 74.46 74.60 74.60 74.57 74.60
Table 10: Results of the best model configurations with different approaches to merge segments during the prediction step. Maximum 4 segments
@@
avg 74.31 74.39 74.33 74.24 73.88
Table 11: Statistics of long documents and distant coreference relations across the datasets
@@ 6.2 Cross-lingual Transfer
Table 12: Zero-shot cross-lingual transfer results. The table compares four settings. A joined model trained for all datasets, a joined model fine-tuned on each dataset separately afterwards, zero-shot setting for each dataset, and zero-shot setting for each language.
@@
Table 12 shows the results of cross-lingual transfer evaluation. We can see that for most of the datasets fine-tuning the model on the specific data helps but the difference is not so significant. For several datasets, the results are even comparable. For small datasets, the fine-tuned models achieve much better results than the joined model. Note that we do not use any weighing of particular datasets in the joined collection so small datasets have a marginal effect on the training. The results of dataset zero-shot and language zero-shot are mostly very predictable. For the languages that have more datasets in CorefUD collection, the drop of zero-shot compared to the joined model was smaller than for the rest of the datasets. When we discard all the datasets for each language, the results are mostly significantly worse compared to discarding just single dataset. This behavior is expected. Several exceptions should be analyzed more. The strangest results are on the Turkish dataset. Here the zero-shot model achieves the best performance. Together with the fact that the numbers are very low for this dataset, it suggests that there is some noise in the dataset annotation. If we train the model on the training part of the Turkish dataset the noise is present on two sides (training and testing). If we train a zero-shot model we eliminate one part.
@@ 6.3 Final Results
Table 13 presents the final results of the best-performing models on a test set. For comparison, we use the best model of CRAC 2023 Multilingual coreference resolution shared task from [37] but not the best submission. The best model is much larger than ours, and it uses ensembling, so as the main model for comparison, we use their RemBERT version without ensembling, which has a comparable size to our model.
From the table, we can see that our model outperformed CorPipe on most of the datasets. For some datasets (mostly the ones without singletons annotated), it outperformed even the large CorPipe model, which has approximately 3 times more trainable parameters. The only dataset where CorPipe outperformed our model by a large margin is the Polish one, which we consider an anomaly because, for this dataset, CorPipe results on the test set are much better than for the dev set. On the dev set, we achieve similar results. The same happened in the opposite way for the Russian dataset. Surprisingly, we even outperformed the large CorPipe model in average score, but this is caused mainly by a large margin on Turkish and German-parcor.
@@ 7 Conclusion
In this paper, we explored and evaluated various approaches to multilingual coreference resolution using the CorefUD 1.1 dataset. Our experiments revealed that monolingual models typically outperform cross-lingual models, especially for languages with datasets that are distinct in their characteristics. However, joint training across languages can still provide benefits for most of the datasets where sufficient cross-linguistic similarities exist.
We proposed several extensions to enhance the baseline models, including cross-lingual training, Span2Head modeling, syntactic information integration, headword mention representation, and long-context prediction. Among these, the heads-only model and singleton modeling showed the most consistent improvements across different datasets, demonstrating the importance of targeted adaptations for coreference resolution tasks. For several datasets, long-context prediction brings also a significant improvement. Additionally, our zero-shot cross-lingual experiments provided insights into the challenges and opportunities of cross-lingual transfer, with the Turkish dataset results highlighting potential issues related to noise in the training data.
@@ Acknowledgement
Computational resources were supplied by the project “e-Infrastruktura CZ” (e-INFRA LM2018140) provided within the program Projects of Large Research, Development and Innovations Infrastructures. This work has been supported by Grant No. SGS-2022-016 Advanced methods of data processing and analysis.